a box contains 9 new light bulbs and 6 used light bulbs. Each light bulb is the same size and shape meredith will randomly select 2 light bulbs from the box without replacement what is the probability meredith will select a new light bulb and then a used light bulb?
Answer by htmentor(1343)

The probability of selecting a new bulb with the 1st pick = P(new) = 9/15
Then the probability of selecting a used bulb with the 2nd pick = P(old) = 6/14
The probability of new followed by old = (9/15)(6/14) = 9/35